The Swift Meal has two restaurants that are open 24 hours a day. Fixed costs for the two restaurants together total $459,000 per year. Service varies from a cup of coffee to full meals. The average sales check per customer is $8.50. The average cost of food and other variable costs for each customer is $3.40. The income tax rate is 30​%. Target net income is $107,100.

Requirements

1. Compute the revenues needed to earn the target net income.

2. How many customers are needed to break​ even? To earn net income of $107,100​?

3. Compute net income if the number of customers is 140,000.
Requirement 1. Compute the revenues needed to earn the target net income.

Revenues

-

Variable costs

-

Fixed costs

=

Target net income

/

1 – Tax rate

The revenues needed to earn the target net income is $

.

Requirement 2. How many customers are needed to break​ even? To earn net income of $107,100​?

Determine the formula that is used to compute how many customers are needed to break​ even, then compute the number of customers needed.

Fixed costs

/

Contribution margin per customer

=

Customers needed to break even

Determine the formula that is used to compute how many customers are needed to earn net income of $107,100 , then compute the number of customers needed.

Total revenues

/

Sales check per customer

=

Customers needed to earn $107,100

Requirement 3. Compute net income if the number of customers is 140,000.

Determine the formula that is used to compute the net income when the number of customers is 140,000​,

then compute the net income.

Change in net income

+

Original target net income

=

Net income
Answer #1
1 Gross profit ratio (8.5-3.4)/8.5
60%
revenues needed to earn the target net income = (Fixed Cost + (Target income/(1-tax rate))/GP ratio
revenues needed to earn the target net income = (459000+(107100/(1-0.3))/60%
revenues needed to earn the target net income = (459000+153000)/60%
Revenues needed to earn the target net income = $1,020,000
2 Customer needed for breakeven = Fixed Cost/contribution per customer
Customer needed for breakeven = 459000/(8.5-3.4)
Customer needed for breakeven =             90,000
Customer needed for target income of 107,100 = (Fixed Cost + (Target income/(1-tax rate))/Contribution per customer
Customer needed for target income of 107,100 = (459000+(107100/(1-0.3))/5.1
Customer needed for target income of 107,100 = (459000+153000)/5.1
Customer needed for target income of 107,100 =           120,000
3 Number of customer 140,000
Contribution per customer $5.1
Total contribution $714,000
Less: Fixed Cost $459,000
Income before tax $255,000
Tax @30% $76,500
Net Income $178,500